OpenStudy (anonymous):

pass through(-2,1),perpendicular to y=4x-11

OpenStudy (anonymous):

44

OpenStudy (anonymous):

y=44

OpenStudy (xapproachesinfinity):

I'm passing through that point write! do you need something

OpenStudy (anonymous):

y = -1/4 x + 1/2

OpenStudy (xapproachesinfinity):

right now*

OpenStudy (anonymous):

im only in 6th

OpenStudy (xapproachesinfinity):

perpendicular to that line means the two slope are opposite reciprocal of one an another which also means the product of those two slope is -1 m1xm2=-1 if m1=a then m2=-1/a hence m1xm2=-1